# What to align in multimodal contrastive learning?

### v1

@@ 7 Conclusions
Multisensory integration is at the core of human perception, allowing us to build coherent representations of our environment. In this paper, we introduce CoMM, a contrastive multimodal method that enables the integration of multiple modalities in a single multimodal representation space. Unlike existing multimodal contrastive models, CoMM is designed to learn multimodal interactions beyond redundancy, through PID theory. Our controlled experiments on the bimodal Trifeature dataset demonstrate that CoMM successfully learns redundant, unique and synergistic information. In real-life multimodal datasets from Multibench with two and three input modalities, CoMM outperforms existing methods by large margins in almost every case, showing the efficiency and versatility of CoMM to handle data across diverse domains (robotics, healthcare, affective computing, multimedia) and structures (time series, image, text, audio, tabular).
This work offers large avenues for future research on multimodal representation learning, in particular for crafting label-preserving multimodal augmentations not limited to unimodal augmentations. Limitations and perspectives for future research are further discussed in Appendix A. Overall, the simplicity and versatility of CoMM’s design make it a good candidate to learn deep representations of several modalities across domains. It offers promises in better solving real-world problems ranging from neuroscience (Preti & Van De Ville, 2019) and medical imaging (Boehm et al., 2022) to remote sensing (Gómez-Chova et al., 2015).
@@ B.1 Encoder architectures by dataset
  • Trifeature is composed of two visual modalities so we use AlexNet encoder for both modalities with $512$-d embedding space. For CoMM, we remove the last average pooling layer and we apply a linear patch embedding layer (Dosovitskiy et al., 2021) to the $6\times 6$ feature maps as latent converter. We then add fixed 2D sine-cosine positional embeddings (Dosovitskiy et al., 2021). For Cross, we use a linear projector for each encoder ($512$-d output space) and we optimize the CLIP loss (Radford et al., 2021). For Cross+Self, we apply a 3-layers MLP projection head to each encoder to optimize the SSL objectives as in SLIP (Mu et al., 2022) ($1024$-d hidden layers and $256$-d output space) and we use the same projectors and CLIP objective as for Cross.
@@
  • MOSI, UR-FUNNY and MUsTARD contain visual and textual modalities extracted from videos. Similarly to FactorCL (Liang et al., 2023b), we use a 5-head Transformer with 5 layers for each modality with a $40$-d embedding space. We do not use latent converter in this case.
@@
  • Vision&Touch has visual, force-torque and robot proprioception modalities available. For the binary contact prediction task, we only use visual and proprioception data for the experiments with two modalities and we encode images with a ResNet18 (He et al., 2016) ($512$-d output space) and proprioception data with a 5-layer MLP ($512$-d output space), as in the original paper (Lee et al., 2020). Force-torque data are encoded with a 5-layer causal convolutions network (512-d output space) in the experiments with three modalities. For CoMM, we remove the last average pooling layer and apply a patch embedding layer (Dosovitskiy et al., 2021) to the $4\times 4$ feature maps as latent converter. We consider the 1D feature vector of proprioception data as a 1-length sequence in the fusion module. For the end-effector regression task, we use visual and force-torque modalities. We also use ResNet18 as image encoder ($128$-d output space) and a 5-layer causal convolutions network ($128$-d output space) as force encoder (Lee et al., 2020). For CoMM, we add the same latent converter for images as in the previous task and we add a feature tokenizer (Gorishniy et al., 2021) for force-torque embeddings.
@@ B.4 Experimental settings
For all downstream classification tasks, we use SupCon loss (Khosla et al., 2020) to fine-tune CoMM (with no additional parameters). In the case of multi-label classification with MM-IMDb, we use a linear head on top of CoMM and we optimize a cross-entropy loss for each label. For the regression task on Vision&Touch, we also use a linear head and we optimize the MSE loss. We systematically use early-stopping according to the validation accuracy in order to prevent over-fitting on the downstream tasks.
B.5 Additional details for LLaVA-NeXT evaluation
@@ C.1 Data augmentation strategy for time-series
We tested these augmentations along with their composition on four datasets from MultiBench (Liang et al., 2021) and we plot the results in Fig. 6. Overall, we find that composing Gaussian noise and random drop results in best performances across datasets and tasks. This is our default policy in our main experiments. Our proposed multi-drop and multi-crop augmentations can provide better results in some cases (for MIMIC and MuSTARD) but we select the same default augmentations for consistency across all datasets.
@@ C.3 Design choices on Trifeatures experiments
Authors in (Hermann & Lampinen, 2020) experimented with Trifeatures on AlexNet and ResNet-50. Both architectures showed comparable results. Therefore, we chose to use AlexNet in our experiments. For completeness, we ran the same experiments on ResNet-50, which show that CoMM is the only model to learn all interactions, regardless of the architecture.
@@
Table 7: Linear probing accuracy of redundancy (shape), uniqueness (texture) and synergy (color and texture) on bimodal Trifeature dataset, with different backbone encoders. These results are complementary to Footnote 4.
@@ Appendix D Pseudo-code
Algorithm 1 presents the pseudo-code for CoMM’s training. It is written in the general case when we have $n$ modalities. It is complementary to Fig. 3 (main paper), which depicts the case $n=2$. CoMM’s official implementation will be released upon acceptance of this work.
@@ E.2 MultiBench
All the following datasets are pre-processed as described in Liang et al. (2021).
  • MIMIC (Medical Information Mart for Intensive Care(Johnson et al., 2016) is a dataset comprising de-indentified clinical data related to patiens admitted to critical care units at a large Boston-area hospital between 2001 and 2012. It contains information about 53 423 hospital admissions, including 38 597 distinct patients. We use the data as provided by MultiBench (Liang et al., 2021), organized as in (Purushotham et al., 2018). There are two data modalities: first, a time series modality, composed by a set of medical measurements of a given patient taken every hour during 24 hours. Each measurement is a vector of size 12 (i.e., including 12 different measured numerical values). Second, a static modality, including medical information about the patient (age, gender, etc.), represented as a vector of size 5 (tabular data). As in (Liang et al., 2023b), in our experiments we address the binary classification task of predicting whether a patient fits any disease in the ICD-9 code in group 7 (460-519). ICD-9 (International Statistical Classification of Diseases and Related Health Problems) codes are used to classify diseases and a variety of symptoms. Almost every health consition can be assigned a unique ICD-9 code group, where each group includes a set of similar diseases.
@@
  • Vision&Touch (Lee et al., 2020) is a robot manipulation multimodal dataset including visual, force and proprioception data for a peg insertion task. The data is collected from a 7-DoF, torque-controlled Franka Panda robot, with a triangle peg attached to its end-effector. Its goal is to insert the peg into a triangle hole situated in a box attached to a table in front of the robot. By running a random policy (the robot takes random actions) and a heuristic policy (the robot attempts peg insertion), 150 trajectories are recorded, each of them consisting of 1 000 timesteps. These trajectories contain RGB images, depth maps, force, end-effector position and velocity. Following MultiBench (Liang et al., 2021), we consider two tasks on this dataset: (i) the binary task of predicting contact or no contact in the next timestep, and (ii) predicting the end-effector possition (measured in MSE).

### v2

@@ 7 Conclusions
Multisensory integration is at the core of human perception, allowing us to build coherent representations of our environment. In this paper, we introduce CoMM, a contrastive multimodal method that enables the integration of multiple modalities in a single multimodal representation space. Unlike existing multimodal contrastive models, CoMM is designed to learn multimodal interactions beyond redundancy, through Partial Information Decomposition theory. Our controlled experiments on the bimodal Trifeature dataset demonstrate that CoMM successfully learns redundant, unique and synergistic information. In real-life multimodal datasets from Multibench with two and three input modalities, CoMM outperforms existing methods by large margins in almost every case, showing the efficiency and versatility of CoMM to handle data across diverse domains (robotics, healthcare, affective computing, multimedia) and data structures (time series, image, text, audio, tabular).
This work opens large avenues for future research on multimodal representation learning, in particular for crafting label-preserving multimodal augmentations not limited to unimodal augmentations. Limitations and perspectives for future research are further discussed in Appendix A. Overall, the simplicity and versatility of CoMM’s design make it a good candidate to learn deep representations of several modalities across domains. It offers promises in better solving real-world problems ranging from neuroscience (Preti & Van De Ville, 2019) and medical imaging (Boehm et al., 2022) to remote sensing (Gómez-Chova et al., 2015).
Acknowledgments
We thank Jonathan Sauder, Valentin Gabeff and Valérie Zermatten for providing helpful feedback on earlier versions of this work. JCN acknowledges the support from EPFL Science Seed Fund. BD acknowledges the support from the PHRT project number 643.
Ethics statement
We acknowledge that there exist potential privacy risks when using human behavior or medical data. However, we minimize these potential risks by using publicly available datasets that have been carefully collected, respecting participants’ consent, de-identifying medical data and anonymizing video data. In this work, we have followed best practices in maintaining the privacy and safety of these datasets.
Reproducibility statement
To ensure the reproducibility of our work, we have used publicly available datasets from MultiBench (Liang et al., 2021) (MIMIC, MOSI, UR-FUNNY, MUsTARD and Vision&Touch); MM-IMDb (Arevalo et al., 2017), and the synthetic Trifeatures dataset (Hermann & Lampinen, 2020). Details about these datasets and pre-processing steps can be found in Appendix E. Appendix B thoroughly describes neural network architectures, data augmentation strategies, and experimental settings used in all our experiments. We include in Appendix D a pseudo-code for implementing CoMM’s training, and we released our code in this Github repository. We also provide details on processing times and model complexity compared to other multimodal models in Appendix F. Finally, all proofs for our theoretical developments can be found in Appendix G.
@@ B.1 Encoder architectures by dataset
  • Trifeature is composed of two visual modalities, so we use an AlexNet encoder for both modalities with a $512$-d embedding space. For CoMM, we remove the last average pooling layer and we apply a linear patch embedding layer (Dosovitskiy et al., 2021) to the $6\times 6$ feature maps as latent converter. We then add fixed 2D sine-cosine positional embeddings (Dosovitskiy et al., 2021). For Cross, we use a linear projector for each encoder ($512$-d output space) and we optimize the CLIP loss (Radford et al., 2021). For Cross+Self, we apply a 3-layers MLP projection head to each encoder to optimize the SSL objectives as in SLIP (Mu et al., 2022) ($1024$-d hidden layers and $256$-d output space) and we use the same projectors and CLIP objective as for Cross.
@@
  • MOSI, UR-FUNNY and MUsTARD contain visual and textual modalities extracted from videos. Similarly to FactorCL (Liang et al., 2023b), we use a 5-head Transformer with 5 layers for each modality with a $40$-d embedding space. We do not use latent converters in this case.
@@
  • Vision&Touch has visual, force-torque and robot proprioception modalities available. For the binary contact prediction task, we only use visual and proprioception data for the experiments with two modalities and we encode images with a ResNet18 (He et al., 2016) ($512$-d output space) and proprioception data with a 5-layer MLP ($512$-d output space), as in the original paper (Lee et al., 2020). In the experiments with 3 modalities, force-torque data are encoded with a 5-layer causal convolutions network (512-d output space). For CoMM, we remove the last average pooling layer and apply a patch embedding layer (Dosovitskiy et al., 2021) to the $4\times 4$ feature maps as latent converter. We consider the 1D feature vector of proprioception data as a 1-length sequence in the fusion module. For the end-effector regression task, we use visual and force-torque modalities. We use ResNet18 as image encoder ($128$-d output space) and a 5-layer causal convolutions network ($128$-d output space) as force encoder (Lee et al., 2020). For CoMM, we add the same latent converter for images as in the previous task and we add a feature tokenizer (Gorishniy et al., 2021) for force-torque embeddings.
@@ B.4 Experimental settings
For all downstream classification tasks, we use the SupCon loss (Khosla et al., 2020) to fine-tune CoMM (with no additional parameters). In the case of multi-label classification with MM-IMDb, we use a linear head on top of CoMM and we optimize a cross-entropy loss for each label. For the regression task on Vision&Touch, we also use a linear head and we optimize the MSE loss. We systematically use early-stopping according to the validation accuracy in order to prevent over-fitting on the downstream tasks.
B.5 Experimental settings on the Bimodal Trifeature dataset
To generate our trifeature dataset, we considered the 1 000 combinations of the three features existing in the original dataset (see Appendix E) and split them into 800 combinations for training and 200 for evaluation. To have more variety in the training set for training, each combination was generated 3 times (the shape and the texture were randomly rotated), obtaining a training split of 2 400 images.
The bimodal Trifeature dataset used in our experiments was built by considering the trifeature dataset twice (as two separate modalities) and building pairs from these two dataset copies. In total, we get 5 760 000 pairs (2 400 $\times$ 2 400) available for training, and 40 000 ($\times 200$) available for evaluation.
To create a controlled environment for evaluation of multimodal interaction learning, we needed to carefully design tasks where the dominant interaction was clearly defined.
  1. 1.
    To measure uniqueness $U_{1}$ (resp. $U_{2}$), given a pair of trifeature images, the task is to predict the texture of the first (resp. the second) image. The task is then a 10-class classification problem and chance level is at 10%.
  2. 2.
    To measure redundancy $R$, given a pair of trifeature images with the same shape (but different color and texture), the task is to predict the shape of the pair (therefore, the redundant information that can be extracted either from the first or the second image). The task is then a 10-class classification problem and chance level is at 10%.
  3. 3.
    To measure synergy $S$, the definition of the task was more subtle as it should require information from both modalities simultaneously and should not be possible to perform it from one of the images alone. To achieve this, we defined a mapping $\mathcal{M}$ between the ten textures and the ten colors (e.g. stripes=red, dots=green, etc.). Then, given a pair of trifeature images, the task is to predict whether the pair satisfies the mapping or not. The task is then a binary classification problem and chance level is at 50%.
To evaluate these tasks, we built two versions of the bimodal trifeature dataset:
  1. (i)
    For uniqueness and redundancy, we considered 10 000 image pairs (out of the 5 760 000 pairs) for training and 4 096 for testing, that have the same shape (to measure redundancy) and different texture (to measure uniqueness).
  2. (ii)
    For synergy, we considered 10 000 image pairs that respect the mapping $\mathcal{M}$ and used the same test set as before (4 096 image pairs).
B.6 Additional details for LLaVA-NeXT evaluation
@@ C.1 Data augmentation strategy for time-series
We tested these augmentations along with their composition on four datasets from MultiBench (Liang et al., 2021) and we plot the results in Fig. 6. Overall, we find that composing Gaussian noise and random drop results in the best performances across datasets and tasks. This is our default augmentations strategy in our main experiments. Our proposed multi-drop and multi-crop augmentations can provide better results in some cases (for MIMIC and MuSTARD), but we select the same default augmentations for consistency across all datasets.
@@ C.3 Design choices on Trifeatures experiments
Hermann & Lampinen (2020) experimented with Trifeatures using AlexNet and ResNet-50 backbones. Both architectures showed comparable results. Therefore, we chose to use AlexNet in our experiments. For completeness, we ran the same experiments on ResNet-50, which show that CoMM is the only model to learn all interactions, regardless of the architecture.
@@
Table 7: Linear probing accuracy of redundancy (shape), uniqueness (texture) and synergy (color and texture) on bimodal Trifeature dataset, with different backbone encoders. These results are complementary to Fig. 4.
C.4 On the feasibility of Assumption 1
(a) Trifeature dataset (synergy)
(b) MM-IMDb dataset
Figure 7: Downstream task performance as the cropping augmentation strength decreases (for Trifeatures) or when both cropping and masking augmentation strength for image and text respectively decreases. $I_{NCE}(X,X^{\prime})$ is measured using the trained encoder $f_{\theta}$ in Eq. 6. As $I_{NCE}(X,X^{\prime})$ increases, the downstream performance first increases and then decreases, revealing a “sweet spot”. It suggests the existence of an optimal augmentation policy that preserves task-relevant information while removing nuisance features from the multimodal input data, giving more credit to our 1.
1 states the existence of an optimal augmentation policy $\mathcal{T}^{*}$ such that $I(X,t(X))=I(X,Y)$ for any $t\in\mathcal{T}^{*}$, given a task $Y$. Hence, the set of optimal multimodal transformations is task-dependent. However, in practice, we have seen that the augmentations applied in our experiments are general enough to obtain good results in a variety of datasets (7 datasets, 10 tasks, diverse data modalities). These augmentations were chosen according to the findings in unimodal self-supervised learning [2, 3], and by a comprehensive ablation study on the best augmentation strategies applicable to time-series in Section C.1.
Theoretically, the set of multimodal augmentations need to be large enough such that no information about the task $Y$ is lost, but small enough to extract this information only.  Tian et al. (2020b) have referred to this observation as the InfoMin principle. In Section 5, our data augmentation ablation study shows that if the set of augmentations is not large enough, then synergy cannot be learnt (the performance is always random chance). In order to further explore the feasibility of 1 and inspired by the strategy developed by Tian et al. (2020b), we evaluate CoMM by progressively increasing the strength of the data augmentations applied. We use random crop as the augmentation to control in the vision domain, mainly for two reasons: first, it is intuitive that by decreasing the level of cropping (keeping less information about the images), we are destroying the task-relevant information; and second, because it has been empirically demonstrated that cropping is a critical transformation in self-supervised learning for vision Chen et al. (2020a). For the text domain, we use masking as the augmentation to control. More specifically, on Trifeatures we randomly crop the original image in the first modality from 0% up to $x$($x=20$% is the strongest augmentation while $x=60$% is the lightest); and from $x$to 100% ($x=0.05$% is the strongest, $x=15$% the lightest). For MM-IMDb, we also use random crop of the image modality from $x$up to 100% and masking of text with a decreasing probability $x$from 90% (the strongest) to 20% (the lightest).
Our results are shown in Fig. 7. They demonstrate, both in the controlled environment of the bimodal Trifeatures dataset and in the real-world application of MM-IMDb, that the sweet spot of the InfoMin principle can be reached. By gradually increasing the strength of the applied transformations, we enhance model performance by reducing noisy information, up to an optimal point (the sweet spot) where noise is minimized while task-relevant information is preserved. However, applying overly strong augmentations destroys task-relevant information, leading to a degradation in model performance.
@@ Appendix D Pseudo-code
Algorithm 1 presents the pseudo-code for CoMM’s training. It is written in the general case when we have $n$ modalities. It is complementary to Fig. 3 (main paper), which depicts the case for $n=2$. CoMM’s official implementation will be released upon acceptance of this work.
@@ E.2 MultiBench
All the following datasets are pre-processed as described in (Liang et al., 2021).
  • MIMIC (Medical Information Mart for Intensive Care(Johnson et al., 2016) is a dataset comprising de-indentified clinical data related to patients admitted to critical care units at a large Boston-area hospital between 2001 and 2012. It contains information about 53 423 hospital admissions, including 38 597 distinct patients. We use the data as provided by MultiBench (Liang et al., 2021), organized as in (Purushotham et al., 2018). There are two data modalities: first, a time series modality, composed by a set of medical measurements of a given patient taken every hour during 24 hours. Each measurement is a vector of size 12 (i.e., including 12 different measured numerical values). Second, a static modality, including medical information about the patient (age, gender, etc.), represented as a vector of size 5 (tabular data). As in (Liang et al., 2023b), in our experiments we address the binary classification task of predicting whether a patient fits any disease in the ICD-9 code in group 7 (460-519). ICD-9 (International Statistical Classification of Diseases and Related Health Problems) codes are used to classify diseases and a variety of symptoms. Almost every health condition can be assigned a unique ICD-9 code group, where each group includes a set of similar diseases.
@@
  • Vision&Touch (Lee et al., 2020) is a robot manipulation multimodal dataset including visual, force and proprioception data for a peg insertion task. The data is collected from a 7-DoF, torque-controlled Franka Panda robot, with a triangle peg attached to its end-effector. Its goal is to insert the peg into a triangle hole situated in a box attached to a table in front of the robot. By running a random policy (the robot takes random actions) and a heuristic policy (the robot attempts peg insertion), 150 trajectories are recorded, each of them consisting of 1 000 timesteps. These trajectories contain RGB images, depth maps, force, end-effector position and velocity. Following MultiBench (Liang et al., 2021), we consider two tasks on this dataset: (i) the binary task of predicting contact or no contact in the next timestep, and (ii) predicting the end-effector position (measured in MSE).